OpenStudy (anonymous):

A rectangle table is three times as it is wide. If the area is 75ft^2, find the length and width of the table??

OpenStudy (anonymous):

l=3b now use A=l*b

OpenStudy (anonymous):

??? I don't understand

OpenStudy (anonymous):

l stands for length and b for width

OpenStudy (anonymous):

so the length is 3b and how do i get the width?? I am horrible at story problems

OpenStudy (anonymous):

now 75=3b*b Now carry on

OpenStudy (anonymous):

b=5 so l=3b=3*5=15

OpenStudy (anonymous):

I still don't understand. what the length and width will be??